The x and y component of vector p are 7i and 6j. Also x and y component of vector p+q are 10i and 10j respectively. The magnitude of q is?

Grade:12th pass

1 Answers

Vikas TU
6 years ago
Dear Student,
p=7i+6j
p+q=10i+10j
q=(p+q)-p = 10i+10j-7i-6j
=3i+4j
so magnitude of q vector = sqrt(3^2 + 4^2)
=5 units  [Ans]
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)